OpenStudy (anonymous):

f(x)=17x/5lxl (absolute value of x) f(3)=?

OpenStudy (anonymous):

f(x) = 17*x/(5*|x|) f(3) = 17*3/(5*|3|) f(3) = 17*3/(5*3) f(3) = 17/5 = 3.4